OpenStudy (anonymous):

Find the number of positive real zeros of f(x)=-7x^4+49x^3+84x^2

sam (.sam.):

you sure is +84 or -84?

OpenStudy (anonymous):

I think it's +

sam (.sam.):

if + you cant get the roots, i'll just assume its -

OpenStudy (anonymous):

ok :) thanks

sam (.sam.):

\[-7x^4+49x^3-84x^2\] Factor the x^2 out \[x^2(-7x^2+49x-84)\] \[x^2(x-3)(x-4)\] Set it =0 to find the roots, \[(x^2)(x-3)(x-4)=0\] Set each brackets =0, x^2=0 x-3=0 x-4=0 x=0 x=3 x=4

OpenStudy (amistre64):

the descartes rule of signs gives us that there is 1 positive root since ther eis only 1 sign change in the eqs positive position

OpenStudy (amistre64):

its not asking for the roots; just the number of roots that are positive in value

OpenStudy (amistre64):

..and obama ate a dog :/